find the domain of (h/g)(x), given that h(x)=5x/(3x-7) and g(x)=(x^4-1)/(5x-15). State your answer in interval notation.
Answer by stanbon(75887)

find the domain of (h/g)(x), given that h(x)=5x/(3x-7) and g(x)=(x^4-1)/(5x-15). State your answer in interval notation.
-----------------------------------------------
(h/g)(x) = [h(x)]/[g(x)] = [5x/(3x-7)]/[(x^4-1)/(5x-15)]
= [5x/(3x-7)] * [5(x-3)/(x-1)(x+1)(x^2+1)]
---------------
I have factored as far can be done; there are no
factors common to numerator and denominator so no
further simplifying is possible.
==================
Cheers,
Stan H.